# Predicting the Distribution of Treatment Effects: A Covariate-Adjustment Approach

### v1

@@ Abstract
Important questions for impact evaluation require knowledge not only of average effects, but of the distribution of treatment effects. What proportion of people are harmed? Does a policy help many by a little? Or a few by a lot? The inability to observe individual counterfactuals makes these empirical questions challenging. I propose an approach to inference on points of the distribution of treatment effects by incorporating predicted counterfactuals through covariate adjustment. I show that finite-sample inference is valid under weak assumptions, for example when data come from a Randomized Controlled Trial (RCT), and that large-sample inference is asymptotically exact under suitable conditions. Finally, I revisit five RCTs in microcredit where average effects are not statistically significant and find evidence of both positive and negative treatment effects in household income. On average across studies, at least $13.6\%$ of households benefited and $12.5\%$ were negatively affected.
@@ 1 Introduction
for any value of $\delta$, by leveraging covariate-adjustment and predictive models. The setting assumes a binary treatment $D$ and that the researcher has access to a representative sample of $Y(1)$ and $Y(0)$, the potential outcomes under treatment and no treatment respectively, and pre-treatment covariates $X$. This is often the case, for example, in Randomized Controlled Trials (RCTs). Alternatively, one can rely on the selection on observables assumption with a known propensity score (Online Appendix I). I denote $\theta$ by Distributional Treatment Effect (DTE), and the method by CAIDE (Covariate Adjustment for Inference on Distributional Effects). All results are pointwise in $\delta$.
@@
Relative to previous literature, my contribution is to provide a new characterization of the identified set of $\theta$, from which I develop two novel methods for inference. The first is valid in finite samples and requires arguably weak conditions. For example, the conditions are satisfied if the data come from an RCT. To my knowledge, this is the first finite-sample inference result on the DTE in the presence of covariates. The second inference method is asymptotically valid and more powerful than the finite-sample approach. Both inference methods rely on estimating a covariate-adjustment term, and I show that they are robust to its misspecification. If the covariate-adjustment term is correctly specified, then the identified set is sharp. If it is misspecified, the identified set is still valid, although wider. As a consequence, the covariate-adjustment term can be estimated using any method, including machine learning algorithms.
@@
My results can be contrasted with those in Fan and Park (2010), who provide a characterization of the sharp identified set of $\theta$ in the presence of covariates. In their approach to inference, they focus on the case without covariates, since it is difficult to implement inference based on their characterization of the identified set, as discussed in Section 2. In contrast, my characterization of the sharp identified set is more amenable to inference, and it satisfies the robustness property mentioned above.
@@
Asymptotic inference on $\theta$ in the presence of covariates was studied by two recent and independent working papers, Semenova (2023) and Ji et al. (2023). Both papers consider the more general problem of learning averages of intersection bounds, which includes the DTE as a special case. Their approach is based on the dual of an optimization problem, and the confidence intervals of both papers coincide for the DTE. By leveraging the specific structure of the problem of learning the DTE, my main advantage relative to these alternatives is that CAIDE always estimates bounds that are (weakly) narrower, leading to more power under suitable conditions (Appendix A). In practice, this difference can lead to different conclusions in terms of statistical significance, as shown in a Monte Carlo study in Section 5 and in the application in Section 6. Another difference compared to the two papers is that I allow for using the proportion of treated units in-sample instead of the true propensity score, which leads to a further reduction in the length of the confidence intervals (Appendix B). Finally, I show asymptotic exactness under continuous outcomes, while Ji et al. (2023) restrict the support of the outcomes to a finite set in order to derive exactness. Semenova (2023) and Ji et al. (2023) do not provide finite-sample inference.
@@
To understand my approach, consider a scenario where a researcher aims to predict $Y(0)$ using a function of covariates $s(X)$. An example of such a function is the conditional expectation, $s(x)=\mathbb{E}[Y(0)|X=x]$. Ideally, $s(X)$ would be close to $Y(0)$, and the distribution of $Y(1)-s(X)$ would approximate the distribution of treatment effects. However, $s(X)$ is generally an imperfect predictor for two main reasons. First, $Y(0)$ is typically random, meaning that even with many covariates, one can only approximate $Y(0)$ but not predict it perfectly. Second, nonparametric methods, including machine learning algorithms, are often biased and exhibit slow rates of convergence, especially with a high number of covariates. CAIDE is designed to address both of these challenges. In this example, the idea is that even if $s(X)$ is an imperfect predictor, examining the distribution of $Y(0)-s(X)$ (the prediction error) can help correct for the difference between $Y(1)-s(X)$ and the actual treatment effect. Intuitively, if $|Y(0)-s(X)|$ is small (e.g., less than 0.1), the error between $Y(1)-s(X)$ and the treatment effect is also small (at most $0.1$). In general, using the entire distribution of $Y(0)-s(X)$ to correct for prediction error provides more information than simply using an upper bound.
@@
The intuition above does not rely on the form of $s$. In fact, by combining the information from the distributions of both $Y(1)-s(X)$ and $Y(0)-s(X)$, any function $s$ can be used to learn $\theta$. However, when correcting for prediction error, different choices of $s$ can lead to different approximations of the distribution of treatment effects. CAIDE proceeds by estimating an optimal form of $s$, denoted by $s^{*}$ (defined in Theorem 1), which depends on the distributions of both potential outcomes conditional on covariates — that is, it is not the prediction of just one potential outcome. In this case, $s^{*}(X)$ can be interpreted as a covariate-adjustment term, since the same function is subtracted from both $Y(1)$ and $Y(0)$.
@@
The fact that any function $s$ is valid motivates a sample-splitting strategy: one portion of the sample is used to estimate $s^{*}$ with $\hat{s}$ using any machine learning algorithm; in the remaining sample, I calculate a confidence interval for $\theta$ taking $\hat{s}$ as fixed. Once this function is fixed, estimation becomes a low-dimensional problem, similar to Fan and Park (2010), and valid inference can be achieved regardless of the statistical properties of $\hat{s}$. In fact, finite-sample inference on $\theta$ (Theorem 2) is valid for example when data comes from an RCT, without requiring further regularity conditions. Moreover, in a more powerful approach and under suitable conditions, cross-fitting, a more efficient way of sample-splitting, can be used for asymptotic inference (Theorem 4). In both cases, if $\hat{s}$ is a poor estimate of $s^{*}$, the resulting confidence interval is wider but still valid, compared to an estimator $\hat{s}$ that is close to $s^{*}$. Both approaches to inference are compared in Section 5 through a Monte Carlo study.
@@
I apply CAIDE to study heterogeneous effects of microcredit (Section 6). Microcredit has reached more than 170 million borrowers worldwide in 2022 (Convergences, 2023), and has been subject of hundreds of academic papers, including several RCTs. Theory suggests that microborrowing can be either beneficial or detrimental (Banerjee, 2013, Garz et al., 2021), and most empirical evidence indicates that average effects are in general mild if not negligible (Banerjee et al., 2015b, Meager, 2019). Analyses of heterogeneous effects have mostly been limited to Quantile Treatment Effects (QTEs), suggesting mainly positive impacts for some, and rarely negative impacts (Meager, 2022). I revisit five studies which randomized microloans at the individual level and find evidence of heterogeneity, including negative treatment effects for some households. The estimated lower bound on the proportion of individuals who would be better off without the loan varies from $2.8\%$ to $26.0\%$, and the lower bound on the proportion who benefited ranges from $2.9\%$ to $26.1\%$.
@@
The rest of the this paper is organized as follows. Section 1.1 reviews the related literature. Section 2 describes the setup and shows the main identification results of the paper. Section 3 provides finite-sample inference, and Section 4 shows asymptotic properties using cross-fitting. Section 5 presents a Monte Carlo study, and Section 6 the results on the application to microcredit. Finally, Section 7 concludes. Proofs that are not displayed in the main text are deferred to the Appendix.
@@ 1.1 Related Literature
Fan and Park (2010) also characterize sharp bounds on $\theta$ in the presence of pre-treatment covariates $X$, but they do not approach inference based on these bounds. It is difficult to provide inference from their characterization, as discussed in Section 2. In contrast, by incorporating covariates via a covariate-adjustment term, the characterization of the sharp bounds that I provide is amenable to inference.
@@
Kallus (2022) studies the sharp bounds in presence of covariates in the case of binary outcomes, and propose inference under high-level assumptions on the data generating process. Beyond the binary case, inference using the sharp bounds on $\theta$ in the presence of covariates was studied in two recent and independent working papers, Semenova (2023) and Ji et al. (2023). The two papers study the more general problem of averages of intersection bounds, which includes $\theta$ as a particular case. Semenova (2023) provides inference when the conditional distributions of potential outcomes can be estimated uniformly at the $n^{-1/4}$ rate, and potential outcomes are discrete. Ji et al. (2023) uses optimal transport theory to show that solving the dual leads to valid bounds even if conditional distributions are misspecified. In contrast, my approach is based on covariate adjustment. The main differences are that I provide: finite-sample inference; asymptotic inference that is more powerful under suitable conditions (Appendix A); a further reduction in the length of confidence intervals by using the proportion of treated units instead of the propensity score (Appendix B); and asymptotic exactness under continuous outcomes.
@@
This paper also contributes to the literature on using machine learning for estimating heterogeneous treatment effects (Athey and Imbens, 2016; Chernozhukov et al., 2018; Wager and Athey, 2018; Athey et al., 2019; Yadlowsky et al., 2021; Athey et al., 2023; Chernozhukov et al., 2023). Most papers in this literature focus on estimating the point-identified conditional average treatment effect (CATE), and use the estimated CATE to predict individual treatment effects. CAIDE is complementary to these approaches. I focus on the partially-identified DTE $\theta$, which answers a different type of question. For example, if the CATE is zero for all values of $X$, it could still be that many individuals are harmed and many benefit. In this case, the DTE could reveal, for example, the proportion of individuals who are harmed and who benefit.
@@
This paper is also related to the literature on conformal prediction (Vovk et al., 2005; Lei et al., 2018). Lei and Candès (2021) propose a prediction interval to the individual treatment effect, i.e., an interval to cover the random variable $Y(1)-Y(0)$ for a specific individual with a certain probability. In contrast, I focus on the population parameter $\theta=P(Y(1)-Y(0)\leq\delta)$. Also related is the literature on covariate adjustment in the context of estimating the Average Treatment Effect (ATE) in RCTs (e.g., Lin, 2013, Wager et al., 2016, Rothe, 2018, Wu and Gagnon-Bartsch, 2018; review in Van Lancker et al., 2023; comparisons in Kahan et al., 2016, Tackney et al., 2023). Most of this literature focuses on adjusting outcomes for covariates with the goal of improving precision. I use covariate adjustment to improve identification (i.e., to get narrower bounds), and not power, of a different estimator, the DTE.
@@
I add to the development economics literature on evaluating the impacts of microcredit. Metanalyses suggest that average impacts are small if not negligible [^2] Breza and Kinnan (2021) find negative effects of the shutdown of microfinance institutions in Andhra Pradesh, which suggests that microcredit might have large and positive equilibrium effects. (Banerjee et al., 2015b, Meager, 2019), however, there is growing evidence that impact is heterogeneous (Banerjee et al., 2019, Bryan et al., 2021, Meager, 2022, Cueva et al., 2024). Analysis of heterogeneity suggested mostly positive impact in the upper tail of income/profits distributions (Angelucci et al., 2015, Augsburg et al., 2015, Banerjee et al., 2015a, Meager, 2022). Yet, evidence of negative effects has been scarce [^3] Crépon et al. (2015), for example, finds negative quantile treatment effect on profits at the 10th percentile. However, they interpret the evidence as a “reduced form”: “they do not necessarily mean that the impact of getting credit itself has the same heterogeneity”, since take-up is low (randomization is done at village level) and “we do not know where the compliers lie in the distribution of outcomes”. They also suggest the negative effect “might be partially due to long-term investments misclassified as current expenses”, which is supported by a positive estimated QTE on consumption at the 10th percentile.. For example, through a metanalysis of seven studies, Meager (2022) finds “no generalizable negative quantile treatment effects”. I revisit five RCTs in microcredit and find novel evidence of both positive and negative effects.
@@ 2 Setup and Identification via Covariate Adjustment
In this section, I define the setup of the paper and present the identification results of $\theta=P(Y(1)-Y(0)\leq\delta)$ that motivate the confidence intervals proposed in Sections 3 and 4. When learning $\theta$ from a sample and predicted counterfactuals, there are two sources of ambiguity. The first is the usual statistical uncertainty due to random sampling error, that is, the usual uncertainty in estimating a population parameter from a sample. The second is due to the prediction of the counterfactuals. That is, even with a large sample size, the predicted counterfactuals can in general approximate the true counterfactuals, but do so with error. The consequence of this ambiguity is that, even with access to population data, in general one can at best find an interval (or bounds) that contain $\theta$ (see, e.g., Heckman et al., 1997; Manski, 1997).
@@
This section lays out the strategy I use to deal with the latter. The first result is a new identifying equation for $\theta$ in (2.3), that is, an interval, denoted identified set, that contains $\theta$ expressed in terms of the distribution of the observable data. This equation incorporates the information on predicted counterfactuals through a scalar function $s(x)$, which acts as a covariate-adjustment term. The main benefit of (2.3) is that the interval contains $\theta$ for any function $s$. That is, (2.3) might be an outer identified set, but it is always a valid identified set. This is crucial so that the confidence intervals proposed in Sections 3 and 4 are valid regardless of the statistical properties of how $s$ is estimated.
@@
The second result is Theorem 1, which characterizes optimal forms of $s$ as to make the identified set in (2.3) sharp, that is, the narrowest possible interval that has the guarantee to contain $\theta$. As it turns out, in general there is not a single choice of $s$ that makes both lower and upper bounds sharp, but different functions $s$ are considered for each bound. The sharp identified set in Theorem 1 is an alternative characterization to the one provided in Fan and Park (2010). Fan and Park (2010) do not address inference on $\theta$ in the presence of covariates. There are two main challenges for inference based on their identifying equation. The first is that if the conditional distributions of potential outcomes are misspecified, then their characterization leads to invalid bounds, that is, bounds that may not contain $\theta$. This is particularly relevant if the conditional distributions are estimated with bias, which is often the case when using nonparametric methods. The second is that the asymptotic distribution of plug-in estimators is not trivial to characterize, and it depends on the rate of convergence of estimators of the conditional distributions. In contrast, misspecification or a slow rate of convergence of the covariate-adjustment term only affects the width of the bounds in (2.3), and the asymptotic distribution is easier to characterize (Theorem 3). Finally, in subsection 2.1 I discuss cases when $\theta$ is point-identified, which happens when covariates are informative enough to accurately predict the counterfactuals.
@@
Assumption A1(i) of bounded $\mathcal{Y}$ is technical, and is discussed in Remark 2. A1(ii) is the usual unconfoundedness assumption, and it is used for identification of the marginals of $Y(1)$ and $Y(0)$. It often holds for example in Randomized Controlled Trials (RCTs), where $D$ is randomized. Relaxing the assumption of constant propensity score is possible, and I propose two alternatives in Online Appendix I. The first allows for the propensity score to vary across groups of $X$, and the second allows for the propensity score to vary arbitrarily with $X$ as long as $p(x)$ is known.
The starting point for CAIDE is the fact that, for any function $s:\mathcal{X}\rightarrow\mathcal{Y}$, it holds that
@@
Therefore, I propose to bound $\theta$ using the marginal distributions of covariate-adjusted outcomes $Y(1)-s(X)$ and $Y(0)-s(X)$. This is in contrast to Fan and Park (2010), which considers the marginals of $Y(1)$ and $Y(0)$ for inference. A major advantage of this approach is that it is flexible in the covariate adjustment term, in the sense that any $s$ induces bounds that contain $\theta$. This is crucial to ensure that the estimation of $s$, regardless of the method or number of covariates used, leads to valid inference in Sections 3 and 4. The Makarov bounds (Makarov, 1982; Williamson and Downs, 1990), used in Fan and Park (2010), always contain $\theta$:
@@
Because of the equivalence in (2.1), and the validity of the Makarov bounds (2.2), (2.3) contains $\theta$ for any $s$. If covariates are not used, e.g. if $s(x)=0$, then (2.3) is equivalent to (2.2). Different functions $s$ lead to different bounds, with varying length. For specific functions, the induced bounds are sharp, that is, the smallest possible valid bounds given the information from the data. One characterization of these bounds was given by Fan and Park (2010):
@@
The result of Theorem 1 is new. It shows that the sharp bounds on $\theta$ in the presence of covariates can be obtained via covariate-adjustment. That is, for specific functions $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$, the induced bounds in (2.3) are equal to the sharp bounds in (2.5). Note that $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ need not be unique, and they are generally different. Hence, in general there is no single function $s$ that induces both lower and upper bounds to be sharp in (2.3), and in Sections 3 and 4 I propose to estimate both $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$.
@@
Functions $s$ that are different from the optimal $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ can lead to wider bounds, including wider than when $s(x)=0$. For example, consider a scenario where outcomes are constant, such as $Y(1)=1$ and $Y(0)=0$. Then, $\theta=P(Y(1)-Y(0)\leq 0)$ is point-identified and equal to zero. Let $s(X)=0$ for half of the population, and $s(X)=1$ for the other half. Then, $Y(1)-s(X)\in\{0,1\}$ and $Y(0)-s(X)\in\{-1,0\}$, and the bounds induced by $s$ are $[0,1]$.

### v2

@@ Abstract
Important questions for impact evaluation require knowledge not only of average effects, but of the distribution of treatment effects. What proportion of people are harmed? Does a policy help many by a little? Or a few by a lot? The inability to observe individual counterfactuals makes these empirical questions challenging. I propose an approach to inference on points of the distribution of treatment effects by incorporating predicted counterfactuals through covariate adjustment. I show that finite-sample inference is valid under weak assumptions, for example, when data come from a Randomized Controlled Trial (RCT), and that large-sample inference is asymptotically exact under suitable conditions. Finally, I revisit five RCTs in microcredit where average effects are not statistically significant and find evidence of both positive and negative treatment effects in household income. On average across studies, at least $13.6\%$ of households benefited, and $12.5\%$ were negatively affected.
Keywords: Distribution of treatment effects, heterogeneous treatment effects, machine learning, causal inference, microcredit.
JEL classification codes: C12, C14.
@@ 1 Introduction
for any value of $\delta$, by leveraging covariate-adjustment and predictive models. The setting assumes a binary treatment $D$ and that the researcher has access to a representative sample of $Y(1)$ and $Y(0)$, the potential outcomes under treatment and no treatment, respectively, and pre-treatment covariates $X$. For example, this is often the setting in Randomized Controlled Trials (RCTs). Alternatively, one can rely on the selection on observables assumption with a known propensity score (Online Appendix OA.1). I denote $\theta$ by Distributional Treatment Effect (DTE) and the method by CAIDE (Covariate Adjustment for Inference on Distributional Effects). All results are pointwise in $\delta$.
@@
Relative to previous literature, my contribution is to provide a new characterization of the identified set of $\theta$, from which I develop two novel methods for inference. The first is valid in finite samples and requires arguably weak conditions. For example, the conditions are satisfied if the data come from an RCT. To my knowledge, this is the first finite-sample inference result on the DTE that incorporates covariates. The second inference method is asymptotically valid and more powerful than the finite-sample approach. Both inference methods rely on estimating a covariate-adjustment term, and I show they are robust to its misspecification. If the covariate-adjustment term is correctly specified, the identified set is sharp. If it is misspecified, the identified set is still valid, although wider. As a consequence, the covariate-adjustment term can be estimated using any method, including machine learning algorithms.
@@
My results can be contrasted with those in Fan and Park (2010), which characterized the sharp identified set of $\theta$ in the presence of covariates. Their approach to inference focuses on the case without covariates since it is challenging to implement inference based on their characterization of the identified set, as discussed in Section 2. In contrast, my characterization of the sharp identified set is more amenable to inference and satisfies the robustness property mentioned above.
@@
Asymptotic inference on $\theta$ in the presence of covariates was studied by two recent and independent working papers, Semenova (2023) and Ji et al. (2023). Both papers consider the more general problem of learning averages of intersection bounds, which includes the DTE as a particular case. Their approach is based on the dual of an optimization problem, and the confidence intervals of both papers coincide for the DTE. By leveraging the specific structure of the problem of learning the DTE, my main advantage relative to these alternatives is that CAIDE always estimates bounds that are (weakly) narrower, leading to more power under suitable conditions (Appendix A.1). In practice, this difference can lead to different conclusions regarding statistical significance, as shown in a Monte Carlo study in Section 5 and in the application in Section 6. Another difference relative to the two papers is that I allow for using the proportion of treated units in-sample instead of the true propensity score, which further reduces the confidence interval length (Appendix A.2). Finally, I show asymptotic exactness under continuous outcomes, while Ji et al. (2023) restrict the support of the outcomes to a finite set to derive exactness. Semenova (2023) and Ji et al. (2023) do not provide finite-sample inference.
@@
To understand my approach, consider a scenario where a researcher aims to predict $Y(0)$ using a function of covariates $s(X)$. An example of such a function is the conditional expectation, $s(x)=\mathbb{E}[Y(0)|X=x]$. Ideally, $s(X)$ would be close to $Y(0)$, and the distribution of $Y(1)-s(X)$ would approximate the distribution of treatment effects. However, $s(X)$ is generally an imperfect predictor for two main reasons. First, $Y(0)$ is typically random, meaning that even with many covariates, one can only approximate $Y(0)$ but not predict it perfectly. Second, nonparametric methods, including machine learning algorithms, are often biased and exhibit slow convergence rates, especially with many covariates. CAIDE is designed to address both of these challenges. In this example, the idea is that even if $s(X)$ is an imperfect predictor, examining the distribution of $Y(0)-s(X)$ (the prediction error) can help correct for the difference between $Y(1)-s(X)$ and the actual treatment effect. Intuitively, if $|Y(0)-s(X)|$ is small (e.g., less than 0.1), the error between $Y(1)-s(X)$ and the treatment effect is also small (at most $0.1$). In general, correcting for prediction error using the entire distribution of $Y(0)-s(X)$ provides more information than simply using an upper bound.
@@
The intuition above does not rely on the form of $s$. In fact, by combining the information from the distributions of both $Y(1)-s(X)$ and $Y(0)-s(X)$, any function $s$ can be used to learn $\theta$. However, when correcting for prediction error, different choices of $s$ can lead to different approximations of the distribution of treatment effects. CAIDE proceeds by estimating an optimal form of $s$, denoted by $s^{*}$ (defined in Theorem 1), which depends on the distributions of both potential outcomes conditional on covariates — that is, it is not the prediction of just one potential outcome. In this case, $s^{*}(X)$ can be interpreted as a covariate-adjustment term since the same function is subtracted from both $Y(1)$ and $Y(0)$.
@@
The fact that any function $s$ is valid motivates a sample-splitting strategy: one portion of the sample is used to estimate $s^{*}$ with $\hat{s}$ using any machine learning algorithm; in the remaining sample, I calculate a confidence interval for $\theta$ taking $\hat{s}$ as fixed. Once this function is fixed, estimation becomes a low-dimensional problem, similar to Fan and Park (2010), and valid inference can be achieved regardless of the statistical properties of $\hat{s}$. In fact, finite-sample inference on $\theta$ (Theorem 2) is valid when data come from an RCT, without requiring further regularity conditions. Moreover, cross-fitting, a more efficient way of sample-splitting, can be used for asymptotic inference under suitable conditions (Theorem 4). In both cases, if $\hat{s}$ is a poor estimate of $s^{*}$, the resulting confidence interval is wider but still valid compared to an estimator $\hat{s}$ that is close to $s^{*}$. I compare both approaches to inference in Section 5 through a Monte Carlo study.
@@
I apply CAIDE to study heterogeneous effects of microcredit (Section 6). Microcredit has reached more than 170 million borrowers worldwide in 2022 (Convergences, 2023) and has been the subject of hundreds of academic papers, including several RCTs. Theory suggests that microborrowing can be either beneficial or detrimental (Banerjee, 2013, Garz et al., 2021), and most empirical evidence indicates that average effects are in general mild if not negligible (Banerjee et al., 2015b, Meager, 2019). Analyses of heterogeneous effects have mostly been limited to Quantile Treatment Effects (QTEs), suggesting mainly positive impacts for some and rarely negative impacts (Meager, 2022). I revisit five studies that randomized microloans at the individual level and find evidence of heterogeneity, including negative treatment effects for some households. The estimated lower bound on the proportion of individuals who would be better off without the loan varies from $2.8\%$ to $26.0\%$, and the lower bound on the proportion who benefited ranges from $2.9\%$ to $26.1\%$.
@@
The rest of this paper is organized as follows. Section 1.1 reviews the related literature. Section 2 describes the setup and shows the main identification results of the paper. Section 3 provides finite-sample inference, and Section 4 shows asymptotic properties using cross-fitting. Section 5 presents a Monte Carlo study, and Section 6 the results of the application to microcredit. Finally, Section 7 concludes. Proofs not displayed in the main text are deferred to the Appendix.
@@ 1.1 Related Literature
Fan and Park (2010) also characterize sharp bounds on $\theta$ in the presence of pre-treatment covariates $X$, but they do not approach inference based on these bounds. It is challenging to provide inference from their characterization, as discussed in Section 2. In contrast, by incorporating covariates via a covariate-adjustment term, the characterization of the sharp bounds that I provide is amenable to inference.
@@
Kallus (2022) studies the sharp bounds in the presence of covariates in the case of binary outcomes and proposes inference under high-level assumptions on the data-generating process. Beyond the binary case, two recent and independent working papers studied inference using the sharp bounds on $\theta$ in the presence of covariates, Semenova (2023) and Ji et al. (2023). The two papers study the more general problem of averages of intersection bounds, which includes $\theta$ as a particular case. Semenova (2023) provides inference when the conditional distributions of potential outcomes can be estimated uniformly at the $n^{-1/4}$ rate and potential outcomes are discrete. Ji et al. (2023) uses optimal transport theory to show that solving the dual leads to valid bounds even if conditional distributions are misspecified. In contrast, my approach is based on covariate adjustment. The main differences are that I provide finite-sample inference, asymptotic inference that is more powerful under suitable conditions (Appendix A.1), a further reduction in the length of confidence intervals by using the proportion of treated units instead of the propensity score (Appendix A.2), and asymptotic exactness under continuous outcomes.
@@
This paper also contributes to the literature on using machine learning for estimating heterogeneous treatment effects (Athey and Imbens, 2016; Chernozhukov et al., 2018; Wager and Athey, 2018; Athey et al., 2019; Yadlowsky et al., 2021; Athey et al., 2023; Chernozhukov et al., 2023). Most papers in this literature focus on estimating the point-identified conditional average treatment effect (CATE) and use the estimated CATE to predict individual treatment effects. CAIDE is complementary to these approaches. I focus on the partially-identified DTE $\theta$, which answers a different type of question. For example, if the CATE is zero for all values of $X$, many individuals could still be harmed, and many could benefit. In this case, the DTE could reveal, for example, the proportion of individuals who are harmed and who benefit.
@@
This paper is also related to the literature on conformal prediction (Vovk et al., 2005; Lei et al., 2018). Lei and Candès (2021) propose a prediction interval to the individual treatment effect, i.e., an interval to cover the random variable $Y(1)-Y(0)$ for a specific individual with a certain probability. In contrast, I focus on the population parameter $\theta=P(Y(1)-Y(0)\leq\delta)$. Also related is the literature on covariate adjustment in the context of estimating the Average Treatment Effect (ATE) in RCTs (e.g., Lin, 2013, Wager et al., 2016, Rothe, 2018, Wu and Gagnon-Bartsch, 2018; review in Van Lancker et al., 2023; comparisons in Kahan et al., 2016, Tackney et al., 2023). Most of this literature focuses on adjusting outcomes for covariates to improve precision. I use covariate adjustment to improve identification (i.e., to get narrower bounds), not the power, of a different estimator, the DTE.
@@
I add to the development economics literature on evaluating the impacts of microcredit. Metanalyses suggest that average impacts are small if not negligible [^2] Breza and Kinnan (2021) find negative effects of the shutdown of microfinance institutions in Andhra Pradesh, which suggests that microcredit might have large and positive equilibrium effects. (Banerjee et al., 2015b, Meager, 2019), however, there is growing evidence that impact is heterogeneous (Banerjee et al., 2019, Bryan et al., 2021, Meager, 2022, Cueva et al., 2024). Analysis of heterogeneity suggested mostly positive impact in the upper tail of income/profits distributions (Angelucci et al., 2015, Augsburg et al., 2015, Banerjee et al., 2015a, Meager, 2022). Yet, evidence of negative effects has been scarce [^3] Crépon et al. (2015), for example, finds negative quantile treatment effect on profits at the 10th percentile. However, they interpret the evidence as a “reduced form”: “they do not necessarily mean that the impact of getting credit itself has the same heterogeneity”, since take-up is low (randomization was conducted at the village level) and “we do not know where the compliers lie in the distribution of outcomes”. They also suggest the negative effect “might be partially due to long-term investments misclassified as current expenses”, supported by a positive estimated QTE on consumption at the 10th percentile.. For example, through a metanalysis of seven studies, Meager (2022) finds “no generalizable negative quantile treatment effects”. I revisit five RCTs in microcredit and find novel evidence of both positive and negative effects.
@@ 2 Setup and Identification via Covariate Adjustment
In this section, I define the setup of the paper and present the identification results of $\theta=P(Y(1)-Y(0)\leq\delta)$ that motivate the confidence intervals proposed in Sections 3 and 4. There are two sources of ambiguity when learning $\theta$ from a sample and predicted counterfactuals. The first is the usual statistical uncertainty due to random sampling error, the usual uncertainty in estimating a population parameter from a sample. The second is due to the prediction of the counterfactuals. Even with a large sample size, the predicted counterfactuals can generally approximate the true counterfactuals, but they do so with error. The consequence of this ambiguity is that, even with access to population data, in general, one can at best find an interval (or bounds) that contain $\theta$ (see, e.g., Heckman et al., 1997; Manski, 1997).
@@
This section lays out the strategy I use to deal with the latter. The first result is a new identifying equation for $\theta$ in (2.3), that is, an interval, denoted identified set, that contains $\theta$ expressed in terms of the distribution of the observable data. This equation incorporates the information on predicted counterfactuals through a scalar function $s(x)$, which acts as a covariate-adjustment term. The main benefit of (2.3) is that the interval contains $\theta$ for any function $s$. That is, (2.3) might be an outer identified set, but it is always a valid identified set. This property is crucial so that the confidence intervals proposed in Sections 3 and 4 are valid regardless of the statistical properties of how $s$ is estimated.
@@
The second result is Theorem 1, which characterizes optimal forms of $s$ as to make the identified set in (2.3) sharp, that is, the narrowest possible interval that has the guarantee to contain $\theta$. As it turns out, in general, no single choice of $s$ makes both lower and upper bounds sharp; instead, each bound has a different function $s$. The sharp identified set in Theorem 1 is an alternative characterization to the one provided in Fan and Park (2010). Fan and Park (2010) do not address inference on $\theta$ in the presence of covariates. There are two main challenges for inference based on their identifying equation. First, if the conditional distributions of potential outcomes are misspecified, their characterization leads to invalid bounds, which may not contain $\theta$. The issue of misspecification is particularly relevant if the conditional distributions are estimated with bias, which is often the case when using nonparametric methods. The second is that the asymptotic distribution of plug-in estimators is not trivial to characterize, and it depends on the rate of convergence of the conditional distribution estimators. In contrast, misspecification or a slow rate of convergence of the covariate-adjustment term only affects the width of the bounds in (2.3), and the asymptotic distribution is easier to characterize (Theorem 3). Finally, in subsection 2.1 I discuss cases when $\theta$ is point-identified, which happens when covariates are informative enough to predict the counterfactuals accurately.
@@
Assumption A1(i) of bounded $\mathcal{Y}$ is technical, and is discussed in Remark 2. A1(ii) is the usual unconfoundedness assumption, and it is used for identification of the marginals of $Y(1)$ and $Y(0)$. It often holds, for example, in Randomized Controlled Trials (RCTs), where $D$ is randomized. Relaxing the assumption of constant propensity score is possible, and I propose two alternatives in Online Appendix OA.1. The first allows the propensity score to vary across groups of $X$, and the second allows it to vary arbitrarily with $X$ as long as $p(x)$ is known.
The starting point for CAIDE is the fact that for any function $s:\mathcal{X}\rightarrow\mathcal{Y}$, it holds that
@@
Therefore, I propose to bound $\theta$ using the marginal distributions of covariate-adjusted outcomes $Y(1)-s(X)$ and $Y(0)-s(X)$. This strategy is in contrast to Fan and Park (2010), which considered the marginals of $Y(1)$ and $Y(0)$ for inference. A major advantage of my approach is that it is flexible in the covariate adjustment term, in the sense that any $s$ induces bounds that contain $\theta$. Such flexibility is crucial to ensure that the estimation of $s$, regardless of the method or number of covariates used, leads to valid inference in Sections 3 and 4. The Makarov bounds (Makarov, 1982; Williamson and Downs, 1990), used in Fan and Park (2010), always contain $\theta$:
@@
Because of the equivalence in (2.1), and the validity of the Makarov bounds (2.2), (2.3) contains $\theta$ for any $s$. If covariates are not used, e.g. if $s(x)=0$, then (2.3) is equivalent to (2.2). Different functions $s$ lead to different bounds, with varying lengths. For specific functions, the induced bounds are sharp, that is, the smallest possible valid bounds given the information from the data. Fan and Park (2010) derived one characterization for these bounds:
@@
The result of Theorem 1 is new. It shows that the sharp bounds on $\theta$ in the presence of covariates can be obtained via covariate adjustment. That is, for specific functions $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$, the induced bounds in (2.3) are equal to the sharp bounds in (2.5). Note that $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ need not be unique and are generally different. Hence, in general, there is no single function $s$ that induces both lower and upper bounds to be sharp in (2.3), and in Sections 3 and 4 I propose to estimate both $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$.
@@
Functions $s$ that are different from the optimal $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ can lead to wider bounds, including wider than when $s(x)=0$. For example, consider a scenario where outcomes are constant, such as $Y(1)=1$ and $Y(0)=0$. Then, $\theta=P(Y(1)-Y(0)\leq 0)$ is point-identified and equal to zero. Let $s(X)=0$ for half of the population and $s(X)=1$ for the other half. Then, $Y(1)-s(X)\in\{0,1\}$ and $Y(0)-s(X)\in\{-1,0\}$, and the bounds induced by $s$ are $[0,1]$.